### 18 Mar 2026 v2

# Uncertainty Relation for Entropy and Temperature of Gibbs States

Abstract
We derive the quantum Fisher information for entropy estimation in a Gibbs state and show that $F_{S}=1/C_{\!v}$, dual to the temperature Fisher information $F_{T}=C_{\!v}/T^{2}$. Their product $F_{S}\cdot F_{T}=1/T^{2}$ is independent of the Hamiltonian, yielding the universal uncertainty relation $\Delta^{2}S\,\Delta^{2}T\geq T^{2}/n^{2}$ in which all system-specific quantities such as heat capacity, the Hamiltonian, and the number of degrees of freedom cancel identically. This is the metrological expression of the Legendre conjugacy between $S$ and $T$. We identify energy measurement as the optimal protocol for entropy estimation, analyse critical-point scaling where $F_{S}\sim|t|^{\alpha}\to 0$, and connect $F_{S}$ to the Ruppeiner metric in entropy coordinates. The uncertainty relation is shown to hold for all standard thermodynamic conjugate pairs, and we examine the distinguished role of the von Neumann entropy within the Rényi family. Generalisations to the grand canonical and generalised Gibbs ensembles are given.
I Introduction
The Legendre structure of thermodynamics pairs each intensive variable with a conjugate extensive one: temperature $T$ with entropy $S$, pressure $P$ with volume $V$, chemical potential $\mu$ with particle number $N$. Among these, the pair $(T,S)$ occupies a distinguished role. Temperature governs the statistical weight of microstates, $p_{k}\propto e^{-E_{k}/T}$, and temperature estimation from quantum measurements of thermal states has a well-developed metrological theory [2, 17, 4, 11, 6]. The quantum Fisher information (QFI) for temperature is $F_{T}=C_{\!v}/T^{2}$ (we set $k_{\mathrm{B}}=1$ throughout), yielding the Cramér–Rao bound $\mathrm{Var}(\hat{T})\geq T^{2}/(nC_{\!v})$. Yet the dual question, with what precision can entropy itself be inferred from quantum measurements, has not been addressed.
This gap is notable, as entropy is a central quantity not only in theory but also in experiment, where its determination via thermodynamic integration [15] is a key tool in systems such as ultracold quantum gases and quantum dots [3, 7]. Maxwell relations also provide a complementary route to thermometry: high precision in temperature estimation occurs when the entropy changes rapidly with respect to a control parameter [13], and the interplay between thermometric and entropic sensitivity in nanoelectronic devices has been explored through the Fisher information [14].
In this Letter we derive the QFI for entropy estimation, show that it is dual to the temperature QFI, and demonstrate that their product yields a universal uncertainty relation whose form is dictated entirely by the Legendre structure of thermodynamics.
II Entropy Fisher information
Consider $n$ independent copies of a Gibbs state $\rho(\beta)=e^{-\beta H}/Z$ at inverse temperature $\beta=1/T$. Because $\rho(\beta)$ belongs to an exponential family, its eigenvalues $p_{k}=e^{-\beta E_{k}}/Z$ are determined by $\beta$ alone and the energy eigenbasis $\{|k\rangle\}$ is $\beta$-independent. The QFI with respect to $\beta$ is therefore purely classical [2]:
$F_{\beta}=\sum_{k}\frac{(\partial_{\beta}p_{k})^{2}}{p_{k}}=\mathrm{Var}(H)=\frac{C_{\!v}}{\beta^{2}}\,$, (1)
where $C_{\!v}\equiv\beta^{2}\mathrm{Var}(H)$ is the heat capacity. The von Neumann entropy $S=-\mathrm{Tr}(\rho\ln\rho)=\beta\langle H\rangle+\ln Z$ is a smooth, strictly monotone function of $\beta$ ($dS/d\beta=-C_{\!v}/\beta<0$ for any thermodynamically stable system with $C_{\!v}>0$). It therefore provides a valid reparametrisation of the one-dimensional Gibbs manifold, and the QFI transforms covariantly as
$F_{S}=\left(\frac{d\beta}{dS}\right)^{\!2}F_{\beta}=\frac{1}{C_{\!v}}\,.$ (2)
The Cramér–Rao bound gives
$\mathrm{Var}(\hat{S})\geq\dfrac{1}{nF_{S}}=\dfrac{C_{\!v}}{n}.$ (3)
This bound is saturated asymptotically by a projective energy measurement (see Appendix). Equation (2) indicates a duality: the heat capacity governs the precision of $S$- and $T$-estimation inversely. Having a large $C_{\!v}$ allows for precise temperature estimation ($F_{T}\propto C_{\!v}$) as small temperature differences produce large, distinguishable shifts in the energy distribution. Conversely, large $C_{\!v}$ makes entropy hard to estimate ($F_{S}\propto 1/C_{\!v}$) because many distinct entropy values map to nearly identical energy statistics.
III Universal uncertainty relation
The product of the two Fisher informations is
$F_{S}\cdot F_{T}=\frac{1}{C_{\!v}}\cdot\frac{C_{\!v}}{T^{2}}=\frac{1}{T^{2}}\,$, (4)
in which $C_{\!v}$ cancels identically. Applying the Cramér–Rao bounds to both variables for $n$ independent copies gives
$\Delta^{2}S\;\Delta^{2}T\;\geq\;\frac{T^{2}}{n^{2}}\,$, (5)
[…]
Crucially, this bound is saturatable: because projective energy measurement simultaneously saturates both individual Cramér–Rao bounds, the product $\Delta^{2}S\,\Delta^{2}T=T^{2}/n^{2}$ is achievable, distinguishing Eq. (5) from a generic (and potentially vacuous) product of independent lower bounds. Equation (5) is the metrological signature of thermodynamic conjugacy: for any pair of Legendre-conjugate variables, an analogous relation holds in which all response functions cancel.
The relation is structurally parallel to the Heisenberg uncertainty principle $\Delta q\,\Delta p\geq\hbar/2$, with $T/n$ playing the role of $\hbar/2$. The analogy is suggestive but imperfect: the Heisenberg relation is state-independent and follows from the canonical commutation relation, whereas Eq. (5) is statistical ($\propto 1/n^{2}$) and restricted to Gibbs states. What distinguishes it from a generic product of Cramér–Rao bounds is the complete cancellation of all system-specific quantities.
IV Two-level system
For a two-level system with energy gap $\Delta$, the heat capacity is $C_{\!v}=(\beta\Delta)^{2}e^{\beta\Delta}/(1+e^{\beta\Delta})^{2}$, giving
$F_{S}=\frac{(1+e^{\beta\Delta})^{2}}{(\beta\Delta)^{2}\,e^{\beta\Delta}}\,.$ (6)
At high temperature ($\beta\Delta\to 0$), $C_{\!v}\to 0$ and $F_{S}\to\infty$: the state approaches the maximally mixed state ($S=\ln 2$), which is sharply localised in entropy space. At low temperature ($\beta\Delta\to\infty$), $C_{\!v}\to 0$ exponentially and $F_{S}\to\infty$ again: the nearly pure ground state has entropy close to zero and is equally well localised. The minimum of $F_{S}$—the point of greatest difficulty for entropy estimation—occurs at intermediate temperature where $C_{\!v}$ is largest. At every temperature, $F_{S}\cdot F_{T}=1/T^{2}$ holds exactly.
Figure 1: Fisher information for entropy and temperature estimation in (a) a two-level system with energy gap $\Delta$ and (b) a quantum harmonic oscillator with frequency $\omega$. Solid curves show $F_{S}=1/C_{\!v}$ (blue) and $F_{T}=C_{\!v}/T^{2}$ (red); the dashed curve is their product $F_{S}\cdot F_{T}=1/T^{2}$.
V Quantum harmonic oscillator
For a quantum harmonic oscillator at frequency $\omega$, the heat capacity is
$C_{\!v}=\frac{(\beta\hbar\omega)^{2}\,e^{\beta\hbar\omega}}{(e^{\beta\hbar\omega}-1)^{2}}\,$, (7)
so $F_{S}=(e^{\beta\hbar\omega}-1)^{2}/[(\beta\hbar\omega)^{2}\,e^{\beta\hbar\omega}]$. Unlike the two-level system, $C_{\!v}$ is monotonically increasing: the spectrum is unbounded and there is no Schottky anomaly, so the metrological structure is qualitatively different. As $T$ increases, entropy estimation steadily approaches a constant while temperature estimation steadily worsens.
The condition $F_{S}=F_{T}$ reduces to $C_{\!v}=T$ , which has two solutions: $T_{1}\approx 0.22\,\hbar\omega$ and $T_{2}\approx 0.90\,\hbar\omega$. Below ($T\ll T_{1}$), the nearly pure ground state is sharply localised in entropy space and $F_{S}$ dominates; The regime ($T_{1}<T<T_{2}$), $F_{T}>F_{S}$, favours thermometry over entropy estimation. In the classical regime ($T\gg T_{2}$), equipartition sets in and the two informations assume their classical asymptotic values. In the classical limit, a system of $f$ quadratic degrees of freedom has $C_{\!v}\to f/2$, yielding:
$F_{S}\;\to\;\frac{2}{f}\,,\qquad F_{T}\;\to\;\frac{f}{2T^{2}}\,.$ (8)
Each additional degree of freedom dilutes the per-copy information about entropy, making $S$ harder to estimate: the energy distribution broadens as $\mathrm{Var}(H)=fT^{2}/2$, so a wider range of entropy values produces statistically similar measurements. The approach to this classical floor is not smooth but reflects the discrete activation of internal modes. As each degree of freedom unfreezes with increasing temperature, $C_{\!v}$ rises in a characteristic staircase; dually, $F_{S}$ descends in discrete steps, providing a metrological signature of degree-of-freedom activation complementary to the familiar staircase in $C_{\!v}$ (see Appendix).
As for all Gibbs states, the product $F_{S}\cdot F_{T}=1/T^{2}$ holds at every temperature, independent of $f$. The number of active degrees of freedom governs how this fixed metrological budget is partitioned between entropy and temperature precision, but cannot alter the budget itself.
VI Critical-point scaling
At a second-order phase transition with specific-heat exponent $\alpha$, $C_{\!v}\sim|t|^{-\alpha}$ where $t=(T-T_{c})/T_{c}$. From Eq. (2),
$F_{S}\sim|t|^{\alpha}\;\to\;0\quad\text{as}\quad t\to 0\,.$ (9)
The Cramér–Rao bound $\mathrm{Var}(\hat{S})\geq C_{\!v}/n$ diverges: near criticality, the proliferation of low-energy excitations renders the energy distributions at neighbouring entropy values statistically indistinguishable. For $\alpha=0$ (e.g. the 2D Ising model [16]), the vanishing is logarithmic: $F_{S}\sim 1/\ln|t|^{-1}$. For a finite system of linear extent $L$, standard scaling gives $C_{\!v}(t,L)\sim L^{\alpha/\nu}\,\tilde{c}(tL^{1/\nu})$, so
$F_{S}(t,L)\sim L^{-\alpha/\nu}\,\tilde{F_{S}}(tL^{1/\nu})\,.$ (10)
At $t=0$ the entropy Fisher information thus decays as a power of system size, providing a direct metrological route to the ratio $\alpha/\nu$: the slope of $\ln F_{S}$ versus $\ln L$ at criticality equals $\alpha/\nu$. This behaviour has a dual interpretation. The thermodynamic length in entropy coordinates [20, 5],
$\mathcal{L}=\int\frac{dS}{\sqrt{C_{\!v}}}\,$, (11)
also vanishes as $C_{\!v}\to\infty$: entropy-changing protocols become quasi-statically cheap near the critical point. The divergence of $C_{\!v}$ therefore simultaneously makes entropy hardest to estimate ($F_{S}\to 0$) and cheapest to change ($\mathcal{L}\to 0$), a complementarity between metrological distinguishability and thermodynamic cost at criticality.
At first-order transitions, the entropy is discontinuous and the Gibbs family breaks down; the QFI framework does not apply across the coexistence region.
VII Relation to thermodynamic geometry
Ruppeiner [18, 19] defined a Riemannian metric on equilibrium states via $g^{R}_{UU}=-\partial^{2}S/\partial U^{2}=1/(T^{2}C_{\!v})$. Janyszek and Mrugała [9] showed that this coincides with the Fisher information metric for Gibbs distributions. Under the coordinate change dU=TdSdU=T\,dS, the metric transforms to $g^{R}_{SS}=1/C_{\!v}=F_{S}$. Thus $F_{S}$ is precisely the Ruppeiner metric component in entropy coordinates, an identification that is implicit in the information-geometry literature but whose estimation-theoretic content has not previously been exploited.
The QFI framing contributes three elements beyond the geometric picture: (i) the bound $\mathrm{Var}(\hat{S})\geq C_{\!v}/n$ is operationally achievable, not merely a measure of statistical distance; (ii) the optimal measurement (projective energy measurement) is identified explicitly; and (iii) the uncertainty relation (5) emerges from combining $F_{S}$ with $F_{T}$, and has no counterpart in the existing thermodynamic geometry literature.
VIII Other conjugate pairs
Our main argument extends to every Legendre-conjugate pair. In each case the relevant Gibbs state is an exponential family in the intensive parameter, the QFI for that parameter equals $\beta^{2}$ times the variance of its conjugate operator, and the Fisher information for the extensive variable follows by reparametrisation. The conjugate susceptibility cancels in the product.
[…]
$F_{V}\cdot F_{P}=\frac{1}{T^{2}}\,$, (12)
with the isothermal compressibility $\kappa_{T}$ cancelling identically from the product (see Appendix for the full derivation). The Cramér–Rao bound gives $\Delta^{2}V\;\Delta^{2}P\geq T^{2}/n^{2}$.
[…]
$F_{N}\cdot F_{\mu}=\frac{1}{T^{2}}\,,\qquad\Delta^{2}N\;\Delta^{2}\mu\;\geq\;\frac{T^{2}}{n^{2}}\,$, (13)
with the particle-number susceptibility $(\partial N/\partial\mu)_{T}$ cancelling identically.
Pair $(\lambda,A)$ $F_{\lambda}$ $F_{A}$ $F_{\lambda}F_{A}$ Bound on $\Delta^{2}\lambda\,\Delta^{2}A$
(T,S)(T,\;S) $C_{\!v}/T^{2}$ $1/C_{\!v}$ $1/T^{2}$ $T^{2}/n^{2}$
[…]
Table 1: Fisher information relations and uncertainty bounds for the three standard thermodynamic conjugate pairs. In every case the conjugate susceptibility cancels identically from the product and the bound $T^{2}/n^{2}$ is universal.
[…]
IX Rényi entropy Fisher information
The von Neumann entropy $S=S_{1}$ is distinguished among the Rényi family $S_{\alpha}=(1{-}\alpha)^{-1}\ln\mathrm{Tr}(\rho^{\alpha})$ by its direct connection to thermodynamic conjugacy. For a Gibbs state, one may show (see Appendix) that $S_{\alpha}(\beta)$ is strictly monotone for every $\alpha>0$, so the reparametrisation argument applies, yielding the Rényi entropy Fisher information
$F_{S}^{\alpha}=\frac{(\alpha{-}1)^{2}\,\mathrm{Var}_{\beta}(H)}{\alpha^{2}\bigl[U(\alpha\beta)-U(\beta)\bigr]^{2}}\equiv\frac{1}{C_{\!v}^{\alpha}}\,$, (14)
where $U(x)\equiv\langle H\rangle_{x}$ is the mean energy at inverse temperature $x$, and $C_{\!v}^{\alpha}$ defines a Rényi heat capacity that reduces to $C_{\!v}$ as $\alpha\to 1$.
The product with the thermometric Fisher information is
$F_{S}^{\alpha}\cdot F_{T}=\frac{(\alpha{-}1)^{2}\,C_{\!v}^{2}}{\alpha^{2}\bigl[U(\alpha\beta)-U(\beta)\bigr]^{2}}\,.$ (15)
For $\alpha\neq 1$, this depends on the Hamiltonian through $U(\alpha\beta)$ and does not reduce to $1/T^{2}$. The universal cancellation of $C_{\!v}$ in $F_{S}\cdot F_{T}$ is therefore special to the von Neumann entropy: it is the unique member of the Rényi family for which $dS/d\beta\propto C_{\!v}$, the same quantity that governs $F_{\beta}$. This singles out $S_{1}$ as the metrologically natural entropy for Gibbs-state estimation.
In the classical limit ($f$ quadratic degrees of freedom), $U(x)=f/(2x)$ and the Rényi Fisher information becomes $F_{S}^{\alpha}\to 2/f$, independent of $\alpha$. This reflects the Gaussian nature of the energy distribution in this regime, where all Rényi entropies are affine functions of $S_{1}$.
X Generalisations
Grand canonical ensemble.—For $\rho(\beta,\mu)=e^{-\beta(H-\mu N)}/\mathcal{Z}$, projecting the two-parameter Fisher matrix onto the entropy direction (see Appendix) yields
$F_{S}^{\mathrm{GCE}}=\frac{1}{\beta^{2}\mathrm{Var}(H-\mu N)}=\frac{1}{C_{\!v}^{(\mu)}}\,$, (16)
where $C_{\!v}^{(\mu)}$ is the heat capacity at constant chemical potential. When $\langle N\rangle$ is additionally held fixed, $F_{S}|_{\langle N\rangle}=1/C_{\!v}^{(N)}$ with $C_{\!v}^{(N)}=\beta^{2}\!\left(\mathrm{Var}(H)-\mathrm{Cov}(H,N)^{2}/\mathrm{Var}(N)\right)$.
Generalised Gibbs ensemble.—For $\rho=e^{-\sum_{k}\lambda_{k}I_{k}}/Z$ with mutually commuting conserved charges $\{I_{k}\}$ and conjugate Lagrange multipliers $\{\lambda_{k}\}$, the entropy $S=\sum_{k}\lambda_{k}\langle I_{k}\rangle+\ln Z$ can be estimated by projecting the multiparameter Fisher matrix $F]_{kl}=\mathrm{Cov}(I_{k},I_{l})$ onto the entropy gradient (see Appendix), giving
$F_{S}^{(\mathrm{GGE})}=\frac{1}{\bm{\lambda}^{T}\!F\,\bm{\lambda}}=\frac{1}{C_{\!v}^{\mathrm{eff}}}\,.$ (17)
In all cases the universal form $F_{S}=1/C_{\!v}^{\mathrm{eff}}$ is preserved and the uncertainty relation (5) holds with $C_{\!v}\to C_{\!v}^{\mathrm{eff}}$. When the conserved charges $\{I_{k}\}$ mutually commute, the SLD operators $L_{\lambda_{k}}\propto(I_{k}-\langle I_{k}\rangle)$ also commute, so the multiparameter SLD bound is simultaneously saturable. In particular, the optimal measurements for all extensive variables—entropy, volume, particle number—are mutually compatible: a single projective measurement in the joint eigenbasis of the charges saturates every scalar Cramér–Rao bound at once. Incompatibility costs arise only for non-Abelian generalised Gibbs ensembles, where the eigenbasis rotates with the parameters.
XI Discussion
The central result of this Letter is that the entropy Fisher information $F_{S}=1/C_{\!v}$ is dual to the thermometric Fisher information $F_{T}=C_{\!v}/T^{2}$, and that their product yields the relation $\Delta^{2}S\,\Delta^{2}T\geq T^{2}/n^{2}$ from which all system-specific quantities cancel by Legendre conjugacy.
Meng and Shi [12] obtained thermodynamic uncertainty relations for parameters encoded in the Hamiltonian, $\hat{H}(\theta)$, with conjugate operator $\hat{O}=\partial_{\theta}\hat{H}$. The $(T,S)$ pair lies outside that framework: $\beta$ parametrises the Gibbs weight rather than the Hamiltonian, and $S$ is a state functional rather than an observable. The universal product $F_{S}\cdot F_{T}=1/T^{2}$ and the resulting cancellation of all system-specific quantities follow instead from the reparametrisation properties of the exponential family.
Experimentally, the bound $\mathrm{Var}(\hat{S})\geq C_{\!v}/n$ is directly relevant to ultracold-atom platforms that determine entropy via thermodynamic integration [15, 21]: it sets the irreducible statistical floor on such measurements given $n$ experimental repetitions. In mesoscopic systems where entropy is accessed via Maxwell relations [3, 7], the bound constrains the precision achievable from charge-sensing measurements. The uncertainty relation (5) further implies that simultaneous high-precision determination of both $S$ and $T$ faces a fundamental trade-off governed only by $T$ and $n$. The connection to fundamental limits of equilibrium metrology [1] suggests broader implications for the design of thermal probes at criticality.
Acknowledgements.
Acknowledgments We would like to thank George Mihailescu, Paolo Abiuso and Daniel Braun for their comments and feedback on the draft. We would like to thank Shubhang Dadhich, Kjartan van Driel, Ned O’Reilly, Marc Nairn, Mathias Bo Svendsen, Parvinder Solanki and particularly Enrique Verduras for fruitful discussions.
Appendix A Reparametrisation validity
The entropy $S(\beta)=\beta\langle H\rangle+\ln Z$ satisfies $dS/d\beta=-C_{\!v}/\beta<0$ for any stable system ($C_{\!v}>0$). It is therefore a smooth, strictly decreasing bijection from $(0,\infty)$ to $(S_{\min},S_{\max})$, establishing a global diffeomorphism between the $\beta$- and $S$-parametrisations of the one-dimensional Gibbs family. Estimating $S$ is operationally equivalent to estimating $\beta$ and applying the deterministic map $\beta(S)$.
The QFI transforms covariantly under any smooth injective reparametrisation [17]: for $\beta\mapsto S(\beta)$,
$F_{S}=\left(\frac{d\beta}{dS}\right)^{\!2}F_{\beta}=\frac{\beta^{2}}{C_{\!v}^{2}}\cdot\frac{C_{\!v}}{\beta^{2}}=\frac{1}{C_{\!v}}\,.$ (18)
This requires: (i) $C_{\!v}>0$ (thermodynamic stability), ensuring injectivity of $S(\beta)$; and (ii) a $\beta$-independent eigenbasis, ensuring $F_{\beta}=\mathrm{Var}(H)=C_{\!v}/\beta^{2}$ is purely classical. Both conditions hold for the canonical Gibbs state.
Appendix B Saturation by energy measurement
The symmetric logarithmic derivative (SLD) of $\rho(\beta)$ with respect to $\beta$ is
$L_{\beta}=\sum_{k}\frac{\partial_{\beta}p_{k}}{p_{k}}|k\rangle\langle k|$ (19)
which is diagonal in the energy eigenbasis. Under the reparametrisation $\beta\mapsto S$, the SLD becomes
$L_{S}=\frac{d\beta}{dS}\,L_{\beta}$ (20)
also diagonal in $\{|k\rangle\}$. The optimal POVM is therefore a projective energy measurement for both $\beta$- and $S$-estimation.
For $n$ independent copies, the sample mean $\bar{E}=n^{-1}\sum_{i}E_{i}$ is a sufficient statistic for $\beta$. The maximum-likelihood estimator $\hat{\beta}_{\mathrm{MLE}}(\bar{E})$ is asymptotically efficient [17], so applying the delta method with $\hat{S}=S(\hat{\beta}_{\mathrm{MLE}})$ gives
$\mathrm{Var}(\hat{S})\;\xrightarrow{n\to\infty}\;\left(\frac{dS}{d\beta}\right)^{\!2}\frac{1}{nF_{\beta}}=\frac{C_{\!v}^{2}/\beta^{2}}{n\,C_{\!v}/\beta^{2}}=\frac{C_{\!v}}{n}\,$, (21)
saturating the Cramér–Rao bound exactly.
Appendix C Grand canonical ensemble
For $\rho(\beta,\mu)=e^{-\beta(H-\mu N)}/\mathcal{Z}$, the Fisher matrix with respect to $(\beta,\mu)$ has entries
$F_{\beta\beta}$ $=\mathrm{Var}(H{-}\mu N)\,$,
$F_{\mu\mu}$ $=\beta^{2}\,\mathrm{Var}(N)\,$, (22)
$F_{\beta\mu}$ $=-\beta\,\mathrm{Cov}(H{-}\mu N,\,N)\,.$
The entropy $S=\beta\langle H-\mu N\rangle+\ln\mathcal{Z}$ depends on both $\beta$ and $\nu=\beta\mu$. Projecting onto the entropy direction via the chain rule gives Eq. (16[10, 8].
Appendix D Generalised Gibbs ensemble
For $\rho=e^{-\sum_{k}\lambda_{k}I_{k}}/Z$ with mutually commuting conserved charges $\{I_{k}\}$, the multiparameter Fisher matrix is $F]_{kl}=\mathrm{Cov}(I_{k},I_{l})$, and the entropy is $S=\sum_{k}\lambda_{k}\langle I_{k}\rangle+\ln Z$. Direct differentiation of $\ln Z$ gives $\partial_{\lambda_{k}}\ln Z=-\langle I_{k}\rangle$, so $\partial_{\lambda_{k}}S=\lambda_{k}[F]_{kk}+\sum_{l\neq k}\lambda_{l}[F]_{kl}$, i.e. $\nabla_{\bm{\lambda}}S=-F\bm{\lambda}$. The scalar Fisher information for $S$ is obtained by the multiparameter chain rule [17],
$F_{S}^{(\mathrm{GGE})}=[(\nabla_{\bm{\lambda}}S)^{T}F^{-1}(\nabla_{\bm{\lambda}}S)]^{-1}=(\bm{\lambda}^{T}F\,\bm{\lambda})^{-1}=\frac{1}{C_{\!v}^{\mathrm{eff}}}\,$, (23)
where $C_{\!v}^{\mathrm{eff}}\equiv(\bm{\lambda}^{T}F\bm{\lambda})^{-1}$ defines the effective heat capacity of the GGE.
Appendix E Metrological equipartition in the classical limit
Consider a collection of oscillators with $f$ total degrees of freedom, $H=\sum_{i=1}^{f}\tfrac{1}{2}\kappa_{i}\xi_{i}^{2}$, where the $\xi_{i}$ are generalised coordinates or momenta with stiffnesses $\kappa_{i}$. In the classical limit $T\gg\max_{i}(\hbar\omega_{i})$, each quadratic degree of freedom contributes $\tfrac{1}{2}k_{\mathrm{B}}T$ to the mean energy and $\tfrac{1}{2}$ to the heat capacity, giving
$C_{\!v}\;\to\;\frac{f}{2}\,,\qquad F_{S}\;\to\;\frac{2}{f}\,,\qquad F_{T}\;\to\;\frac{f}{2T^{2}}\,.$ (24)
The entropy Fisher information per degree of freedom is $F_{S}/f=2/f^{2}$, meaning each additional quadratic mode makes entropy harder to estimate by diluting the per-copy information content. This has a transparent physical origin: with more degrees of freedom the energy distribution broadens ($\mathrm{Var}(H)=fT^{2}/2$), so a wider range of entropy values produces statistically similar energy measurements.
The approach to this classical floor is governed by the quantum corrections. For a single oscillator at frequency $\omega$, expanding $C_{\!v}$ to next order gives $C_{\!v}=1-(\beta\hbar\omega)^{2}/12+\mathcal{O}(\beta^{4})$, so
$F_{S}=1+\frac{(\beta\hbar\omega)^{2}}{12}+\mathcal{O}(\beta^{4})\,.$ (25)
The leading quantum correction increases $F_{S}$: as the temperature drops below the classical regime, the freezing-out of high-energy states begins to constrain the entropy, making it slightly easier to estimate. For a collection of $f$ oscillators with frequencies $\{\omega_{i}\}$,
$F_{S}=\frac{2}{f}\left(1+\frac{\sum_{i}(\beta\hbar\omega_{i})^{2}}{6f}+\mathcal{O}(\beta^{4})\right).$ (26)
The quantum correction is weighted by the mean-square frequency, $\overline{\omega^{2}}=f^{-1}\sum_{i}\omega_{i}^{2}$. Systems with a broad frequency spectrum (e.g. a Debye solid) therefore depart from the equipartition floor at higher temperatures than systems with a narrow spectrum, and the departure is anisotropic in frequency space.
For a three-dimensional monatomic ideal gas ($f=3$ translational modes), $C_{\!v}\to 3/2$ and $F_{S}\to 2/3$. A diatomic gas at temperatures above the rotational but below the vibrational threshold has $f=5$ (three translations, two rotations) giving $F_{S}\to 2/5$; upon activating the vibrational mode, $f\to 7$ and $F_{S}\to 2/7$. Each activation of a new degree of freedom produces a discrete drop in $F_{S}$, providing a metrological signature of the stepwise unfreezing of internal modes that is complementary to the familiar staircase in $C_{\!v}$.
The product relation is of course maintained throughout: $F_{S}\cdot F_{T}=(2/f)(f/2T^{2})=1/T^{2}$, independent of $f$. The number of degrees of freedom governs how the fixed metrological budget $1/T^{2}$ is partitioned between entropy and temperature precision, but cannot alter the budget itself.
Appendix F Other conjugate pairs: detailed derivations
The (P,V)(-P,V) pair
In the isothermal–isobaric ensemble $\rho=e^{-\beta H-\gamma\hat{V}}/Z$ with $\gamma\equiv\beta P$ and $\hat{V}$ the volume operator. At fixed $\beta$, this is an exponential family in $\gamma$ with a $\gamma$-independent eigenbasis, so
[…]
Since $d\gamma/dP=\beta$,
[…]
The fluctuation–response theorem gives $\mathrm{Var}(\hat{V})=T\kappa_{T}V$, where $\kappa_{T}=-(1/V)(\partial V/\partial P)_{T}>0$ is the isothermal compressibility. The equilibrium volume $\langle\hat{V}\rangle(P)$ is strictly monotone (stability requires $\kappa_{T}>0$), so we may reparametrise $P\mapsto V$:
$F_{V}=\left(\frac{dP}{dV}\right)^{\!2}F_{P}=\frac{\beta^{2}\mathrm{Var}(\hat{V})}{(\kappa_{T}V)^{2}}\,.$ (29)
[…]
$F_{V}=\frac{\beta^{2}\mathrm{Var}(\hat{V})}{\beta^{2}\mathrm{Var}(\hat{V})^{2}}=\frac{1}{\mathrm{Var}(\hat{V})}\,$. (30)
The product is
$F_{V}\cdot F_{P}=\frac{1}{\mathrm{Var}(\hat{V})}\cdot\beta^{2}\mathrm{Var}(\hat{V})=\beta^{2}=\frac{1}{T^{2}}\,$, (31)
and the Cramér–Rao bound gives
$\Delta^{2}V\;\Delta^{2}P\;\geq\;\frac{T^{2}}{n^{2}}\,.$ (32)
The compressibility $\kappa_{T}$ cancels identically.
The $(\mu,N)$ pair
In the grand canonical ensemble $\rho=e^{-\beta(H-\mu\hat{N})}/Z$ with $\nu\equiv\beta\mu$, the same argument at fixed $\beta$ gives
[…]
and, using $\mathrm{Var}(\hat{N})=(\partial N/\partial\mu)_{T}/\beta$ and strict monotonicity of $\langle\hat{N}\rangle(\mu)$,
$F_{N}=\frac{1}{\mathrm{Var}(\hat{N})}\,$. (34)
The product $F_{N}\cdot F_{\mu}=\beta^{2}=1/T^{2}$, and the bound is
$\Delta^{2}N\;\Delta^{2}\mu\;\geq\;\frac{T^{2}}{n^{2}}\,.$ (35)
The particle-number susceptibility $(\partial N/\partial\mu)_{T}$ cancels identically.
Appendix G Rényi entropy Fisher information: detailed derivation
For a Gibbs state $\rho(\beta)=e^{-\beta H}/Z(\beta)$, the $\alpha$-power trace evaluates to
$\mathrm{Tr}(\rho^{\alpha})=\frac{Z(\alpha\beta)}{Z(\beta)^{\alpha}}\,$, (36)
so
$S_{\alpha}(\beta)=\frac{1}{1{-}\alpha}\bigl[\ln Z(\alpha\beta)-\alpha\ln Z(\beta)\bigr]\,.$ (37)
Differentiating with respect to $\beta$ and using $d\ln Z(x)/dx=-U(x)$, where $U(x)\equiv\langle H\rangle_{x}$ is the mean energy at inverse temperature $x$, yields
$\frac{dS_{\alpha}}{d\beta}=\frac{\alpha}{\alpha{-}1}\bigl[U(\alpha\beta)-U(\beta)\bigr]\,.$ (38)
For any $\alpha>0$ with $\alpha\neq 1$ and any thermodynamically stable system, $U(x)$ is strictly decreasing (since $dU/dx=-\mathrm{Var}_{x}(H)<0$), so the factor in brackets is non-zero and $S_{\alpha}(\beta)$ is strictly monotone: a valid reparametrisation of the Gibbs manifold for every $\alpha$.
The Rényi entropy Fisher information follows by the same covariant transformation used in the , yielding Eq. (14).